OpenStudy (anonymous):

zero product property: (2x-4)(2x-1)=0 solve the equation

jimthompson5910 (jim_thompson5910):

if (2x-4)(2x-1)=0 then 2x-4=0 or 2x-1=0 solve each equation for x

OpenStudy (anonymous):

that's all i have to do?

jimthompson5910 (jim_thompson5910):

yes you need to solve 2x-4 = 0 for x and solve 2x-1 = 0 for x

OpenStudy (anonymous):

so its 2 and 1/2

jimthompson5910 (jim_thompson5910):

correct the two solutions are x = 2 or x = 1/2

OpenStudy (anonymous):

thanks :))

jimthompson5910 (jim_thompson5910):

np